OpenStudy (anonymous):

Why can 25x^2 + 144 not be rewritten as a group of factors?

OpenStudy (anonymous):

@hartnn do you think you can help me again :)

hartnn (hartnn):

ok, a quadratic expression ax^2+bx+c can be factored if b^2-4ac is a perfect square.

hartnn (hartnn):

so, compare 25x^2+144 with ax^2+bx+c and tell me what u get for a,b,c ?

OpenStudy (anonymous):

ok, 'a' would be 25 and 'c' would be 144 but i don't have a 'b'

OpenStudy (solomonzelman):

b=0

hartnn (hartnn):

yup, 25x^2+144= 25x^2+0x+144

hartnn (hartnn):

so, now whats b^2-4ac = ...?

OpenStudy (anonymous):

so then is would be 0 - 4 (25)(144) = -14,400?

hartnn (hartnn):

is that a perfect square?

OpenStudy (anonymous):

no because it is negative right? so would that be my reason?

hartnn (hartnn):

correct! since the discriminant b^2-4ac for 25x^2+144 is NOT a perfect square, it cannot be factored into groups. :)

OpenStudy (anonymous):

ok :) thank you very much! but do you think you can help me with one more? it is a bit bigger?

hartnn (hartnn):

sure!

OpenStudy (anonymous):

Ok :) so with the same equation, 25x^2 + 144, i have to graph it on a graph. So far i have put it on a graph and it is a positive parabola that is very narrow, other then that i have points the i solved for as well which are (-2, 244), (-1, 169), (0,144), (1,169), and (2, 244). Now i am asked to find the average rate of change and whether it is increasing or decreasing, from left to right. And i also have to justify my observation with calculations. So far i know kind of that you are suppose to have two points (or something around those lines) and find the slope of them. That way i would have to rate of change but i don't know how to solve for the rest...

OpenStudy (anonymous):

i am very sorry how long it is :(

OpenStudy (anonymous):

do you think you can help @hartnn ? i know its a lot and im sorry :(

hartnn (hartnn):

oh, sorry for the late reply! i got busy with these reports... increasing if avg. rate of change (or slope) is positive decreasing if avg. rate of change (or slope) is negative

OpenStudy (anonymous):

That's it? what about the left to right thing? and what about calculations?

OpenStudy (anonymous):

also what would be my two points that i would use?

hartnn (hartnn):

ok, from graph you would observe that it is decreasing from left to right till x=0 and then increasing thereafter. so, first take 2 points on left of x=0 show that slope is negative (calculations part)

hartnn (hartnn):

then take 2 points on right of x=0 show that slope is positive , hence increasing

OpenStudy (anonymous):

Ahh is that all i need?

hartnn (hartnn):

i think yes. that would cover up all questions...

OpenStudy (anonymous):

thank you very very much ! :D

hartnn (hartnn):

observations will be from graph and calculations is of finding slope with the formula. welcome ^_^